OpenStudy (anonymous):

A 340.0−mL buffer solution is 0.140M in HF and 0.140M in NaF. What mass of NaOH could this buffer neutralize before the pH rises above 4.00?

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

@Somy

OpenStudy (aaronq):

Easiest way is to use the Henderson-Hasselbalch equation \(pH=pKa+log\dfrac{[A^-]}{[HA]}\) Find the change to occur is \([A^-]\) consistent to that pH change. Then use \(Molarity=\dfrac{moles_{solute}}{L_{solution}}\) to find the moles used. These moles are equal to the moles of NaOH neutralized. finally, use \(moles=\dfrac{mass}{Molar~mass}\) to find the mass of NaOH.

OpenStudy (aaronq):

the change to occur IN*** \([A^-]\)